Noel's age in 8 years will be 3 times as old as it was 2 years ago. How old is Noel now?
Found 2 solutions by ewatrrr, stanbon:
Answer by ewatrrr(24785)

x Noel's age NOW
...
x+8 = 3(x-2)
x + 8 = 3x - 6
14 = 2x
7 = x, Noel's age NOW
and...checking
15 = 3(5)

Answer by stanbon(75887)

Noel's age in 8 years will be 3 times as old as it was 2 years ago.
How old is Noel now?
-----
Equation:
N+8 = 3(N-2)
N+8 = 3N-6
2N = 14
N = 7 yrs (noel's age now)
------------
Cheers,
Stan H.
------------